find the sum of two even consecutive intergers whose sum is 702
Answer by stanbon(48558)

find the sum of two even consecutive integers whose sum is 702
Let the 1st number be "2x" so it will be even.
The next consecutive even number is "2x+2"
EQUATION:
2x+2x+2 = 702
4x=700
x= 175
2x=350 (1st even number)
2x+2=352 (2nd even number)
Cheers,
Stan H.